# Scaling Towards the Information Boundary of Instruction Sets: The Infinity Instruct Subject Technical Report

### v3

@@
LiDu Hanyu Zhao Yiming Ju Tengfei Pan Li Du1,2 Hanyu Zhao1,2 Yiming Ju1,2 Tengfei Pan
@@
{ldu, hyzhao, ymju, tfpan}@baai.ac.cn
(December 3, 2025)
@@ 2 Introduction
Instruction tuning serves as a cornerstone for unleashing the vast knowledge and reasoning capabilities of large pretrained models (longpre2023flan; ouyang2022training). Consequently, the construction of high-quality instruction datasets has become essential for improving model performance and generalization ability. Several efforts have constructed instruction datasets through manual annotation or automatic synthesis, with the total size of these datasets having reached tens of millions (ahmad2025opencodeinstruct; nayak2024learning). However, current large models still struggle with complex instruction-following tasks and show limitation in “rare” tasks with low frequency (qin2025incentivizing).
A main reason would be that existing works overlooked the distribution of existing data, and thus failed to efficiently enhance the “coverage and “depth of synthesized instructions, and thus limit the model performance in solving rare tasks and difficult tasks (wang2024survey; liu2024best; long2024llms). The “coverage of an instruction dataset refers to the coverage of task types and knowledge categories it covers, while the ‘depth reflects the complexity of instructions within individual tasks. Both two dimensions are essential to enlarge the capability boundaries of large models (shengyu2023instruction). Expanding the coverage improves the model’s ability to generalize across domains, while increasing depth helps enhance complex reasoning ability and compositional generalizability (sinha2024survey; zhong2025survey). However, a key challenge remains: how to design effective, controllable and interpretable instruction generation strategies that continuously expand the instruction space along both dimensions, guiding models to acquire higher-level capabilities (zhang2023survey; qian2022controllable; qin2022cold). Currently, there is a lack of unified methodology and empirical research in this area.
@@
Based on this framework, we construct the Infinity Instruct Subject (InfInstruct-Sub) dataset, which contains $\sim$1.5 million high-quality instructions. The code and dataset are available at https://github.com/BAAI-DIPL/InfinityInstruct-Sub and https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen. Evaluations on multiple benchmarks demonstrate its effectiveness in improving instruction-following ability. Across foundation models, our tuned models outperform their official instruction-tuned counterparts. Data analysis further reveals that, compared with comparable synthetic datasets such as Magpie (magicoder) and UltraChat (ding2023enhancing), our framework could obtain an instruction set with higher coverage and complexity. Moreover, interestingly, we observe a scaling law in the co-occurrence structure of tags of instructions: the co-occurrence frequency of a tag with others follows a negative logarithmic relationship with its own frequency. This indicates a scale-free connection structure between instructions, similar to the topology of the Internet. Such a pattern may offer new insights into the scaling laws of model performance and suggest ways to better understand the internal knowledge structure of datasets, potentially improving training efficiency and effectiveness. This further highlights the importance of understanding the content structure of instruction data, and underscores the value of the high-quality tagging system proposed in this work.
@@ 3 Construction of Infinity Instruct Subject
The construction process of Infinity Instruct Subject (InfInstruct-Sub) is illustrated in Figure 1. The process starts by collecting high-quality seed instructions. We systematically gathered all general-domain instruction datasets both manually created and automatically generated using GPT-4(achiam2023gpt) or ChatGPT available up to March 2024, resulting in a data pool of approximately 7 million samples. The included datasets are listed in Table 1 of the Appendix. Next, we developed an automatic tagging system based on large language models (LLMs) to analyze the distribution of the data pool. From the perspectives of coverage and depth, we selected a set of high-information seed instructions. Then, we applied an evolutionary algorithm to generate over one million new instruction samples by evolving towards greater complexity and difficulty. Building on both the seed and synthesized data, we constructed a model deficiency diagnosis system. This system identifies gaps in model capabilities and guides the targeted synthesis of new data to efficiently address those weaknesses. Finally, we implemented a strict semantic similarity–based data leakage prevention framework to detect and mitigate potential risks of data leakage throughout the construction process.
@@ 3.1 Hierarchical Multilingual Tagging System for Characterizing Instruction Content Distribution
Since LLMs may describe the same knowledge or skill using different expressions (e.g., “math calculation” vs. “mathematical calculation”), we normalize tags based on semantic similarity. Specifically, we obtain embeddings for all tags using BGE (chen2024bge), and identify semantically similar tags whose cosine similarity exceeds an empirical threshold of $\lambda=0.91$. Within each group of similar tags, we retain the one with the highest frequency as the representative. To further refine the tag set, we apply DBSCAN (ester1996density) clustering with parameters $eps=0.47$ and min_samples=2min\_samples=2 to merge closely related tags. These thresholds are chosen based on empirical observations. After normalization and clustering, tags with frequency less than 100 are considered noisy and removed, following prior work (lu2023instag). As a result, a total of 21,378 fine-grained tags are retained.
@@
Given the set of fine-grained tags, we select the top 1,000 most frequent ones, then use GPT-4 to automatically summarize them into a set of domain-level categories, so that we employ each induced category as a domain-level tag. To establish a mapping between fine-grained tags and domain-level tags, we use the prompt shown in Figure 4 along with the categorization criteria illustrated in Figure 1 of the Appendix. Based on this setup, we employ Qwen2.5-72B-Instruct (qwen2.5) to generate the mappings between fine-grained and domain tags.
@@ 4. Undertrained Instructions:
In particular, we use Llama-2-7B-base (touvron2023llama) to compute the loss values for instructions. To estimate the reduction ratio in token-level average cross-entropy loss, we fine-tune Llama-2-7B-base on a balanced instruction subset with 1,000 instructions per domain category. Table 1 presents the sources of the seed data and the number of instructions retained after filtering. By applying the above criteria sequentially, we select approximately 1.2 million informative instructions from the overall data pool.
@@ 3.3 Evolutionary Algorithm for In-Distribution Generalization
1. Metadata-Guided Evolution: In this step, we adopt the widely-adopted Evol-Instruct (xu2023wizardlm) method. By randomly choose one of four dimensions—diversity, i.e., more reasoning steps, concretizing, deepening, and more diversified, to guide each instruction’s evolution using State-of-the-Art advanced large model as synthesizer.
@@
3. Multi-Turn Dialogue Generation: Following (ding2023enhancing), each valid instruction is used to generate 1–4 rounds of dialogue, with the AI assistant simulating different roles.
@@ 3.5 Data Leakage Prevention
Overlap between training and evaluation data can cause models to achieve inflated scores on familiar samples, leading to overestimated performance. We observe that many public instruction datasets include questions from common benchmarks like MT-Bench (zheng2023judging) and AlpacaEval (li2023alpacaeval), which risks misleading evaluation. To address this issue, we use the BGE model to compute semantic similarity between evaluation queries and instructions in the training data. This process involves multi-dimensional comparisons, including semantic understanding and structural features, to ensure accurate and comprehensive matching. Based on the similarity scores, we filter out overlapping entries, effectively preventing data leakage and preserving the reliability of model evaluation.
@@ 4.1 Experimental Setup
We finetune open-source pretrained models Qwen2-7B-base and LLaMA3-8B-base on the InfInstruct-Sub dataset, and compare their performance against their respective official instruction-tuned and alignment-tuned versions. Evaluations are conducted on widely adopted LLM-judging-based benchmarks AlpacaEval 2.0 (li2023alpacaeval) and Arena-Hard-V0.1 (li2024live). The reference model for AlpacaEval 2.0 is GPT-4-1106; on Arena-Hard-V0.1 is GPT-4-0314.
@@ 4.2 Experimental Results
+Self-Instruct (wang2023self) 100k 7.21 4.0
+ShareGPT (zheng2023judging) 112k 9.73 6.5
+Evol Instruct (xu2023wizardlm) 143k 8.52 5.1
+OpenHermes 1 (OpenHermes1) 243K 9.94 4.4
+Tulu V2 Mix (ivison2023camels) 326K 9.91 5.4
+WildChat (zhaowildchat) 652k 14.62 8.7
+OpenHermes 2.5 (OpenHermes2) 1M 12.89 8.2
+GenQA (genqa_dataset) 6.47M 9.05 3.0
+UltraChat (ding2023enhancing) 208k 8.29 3.6
+Magipie (xu2024magpie) 300k 22.66 14.9
@@ 5.1 Dataset Distribution
We further compare the semantic coverage of InfInstruct-Sub with similar instruction datasets, including Alpaca(taori2023stanford), llm-sys(zheng2023lmsys), and Magpie (xu2024magpie). We sample 20,000 instructions from each dataset and apply the embedding and projection method same to abovementioned. As shown in Figure 12, InfInstruct-Sub exhibits a broader semantic spread, demonstrating that our strategy of selecting more low-frequency instructions as seeds effectively increases coverage in the synthesized data.
@@
To evaluate the difficulty of synthesized data, we follow the approach proposed in Magpie (xu2024magpie), using a large language model to assign difficulty scores to instruction samples. The difficulty is rated on a five-point scale: ’very easy’: 0, ’easy’: 1, ’medium’: 2, ’hard’: 3, ’very hard’: 4. This evaluation process is independent of the method we use for selecting high-difficulty seed instructions. Specifically, we employ Qwen-2.5-32B-Instruct as the scorer and randomly sample 1,500 instructions from each of the following datasets: Alpaca-GPT, LLM-Sys, Magpie, and InfInstruct-Sub.
@@ 5.3 Scaling Phenomenon in Label Connectivity Distribution
The presence of this topology highlights the interdependent structure of knowledge in instruction data, which should be considered within the data construction, instruction selection, and model training process. This also highlights the necessity of understanding the distribution and interrelation of instruction contents (zhao2025beyond), and the necessity of a high-quality tagging system, which is the foundation of understanding such knowledge structure.
Moreover, prior work (e.g., (tan2024information)) hypothesizes that the scaling law between model performance and dataset size is originated from the distribution of “latent skills” within the data, if such “latent skills” follow a power-law distribution. Our findings provide empirical support for this hypothesis, offering new insights into the origin of scaling laws between data distribution and model performance. We leave further exploration of this direction to future work, which further emphasizes the importance of the tagging system introduced in this paper.
@@ 6.1 Instruction Data Synthesis
Manually constructed datasets rely on experts to write instructions and responses, such as LIMA (lima2023) and Dolly (ding2023enhancing). While these datasets are high in quality, they are expensive to scale due to the high cost of human annotation.
@@
Semi-automatic approaches, like Self-Instruct(wang2023self), Alpaca(taori2023stanford), and Evol-Instruct(muennighoff2023octopack), expand a small set of human-labeled data using prompt engineering. While they improve scalability, the reliance on handcrafted prompts limits diversity and complexity. Fully automatic methods (e.g., WebInstruct(yue2024mammoth2), backtranslation) extract instruction-like data from web documents with minimal human input, but often lack precise control over coverage and difficulty. Recent studies explore improvements from three angles:
@@
  • Seed Selection and High-Information Filtering: The choice of seed data largely determines the downstream diversity and quality. Traditional methods often use a few high-frequency, simple instructions, leading to limited coverage(wang2023self). Recent work addresses this by leveraging hierarchical tag systems and multi-dimensional metrics to select high-information seeds—i.e., rare, diverse, and complex instructions that better represent undercovered regions of the task space(tu2024resofilter; he2025fine; xie2023data). This ensures that synthesized data can expand both the coverage (domain/task diversity) and depth (instruction complexity) of the dataset.
@@
  • Evolution-Based Instruction Generation: To address the need for more challenging instructions, evolution-based methods iteratively expand seed data using strategies like genetic algorithms (xu2023wizardlm) or model feedback loops(gu2025llm). These approaches increase instruction complexity and reasoning depth, while dynamically adjusting generation based on model performance. They also help avoid redundancy by diversifying synthesis paths.
@@
  • Instruction Synthesis Strategies: Most existing synthesis methods rely on prompting large language models with handcrafted or auto-generated templates(wang2023self; xu2023wizardlm). While effective, this can lead to templated and repetitive outputs. Magpie(xu2024magpie) proposes prompt-free instruction synthesis via autoregressive alignment, which improves fluency and semantic richness. Similarly, Web Reconstruction (jiang2025instruction) leverages real-world text to generate instruction-style data with higher realism and contextual diversity.
@@
Despite these advances, many datasets still underrepresent complex, multi-step, or domain-specific instructions. Few existing works explicitly address the coverage and depth of instruction distribution in a unified manner. This limits model generalization in scenarios requiring deep reasoning, long-range context, or adaptation to unfamiliar domains (yu2024survey; zhang2025out). Moreover, while some studies explore feedback-driven data augmentation or generative instruction synthesis (le2022coderl), most lack a closed-loop framework that jointly optimizes coverage, difficulty, and model adaptation. The interplay between seed selection, model feedback, and evolution remains underexplored.

### v4

@@
LiDu Hanyu Zhao Yiming Ju Chengwei Wu Tengfei Pan Li Du1,2 Hanyu Zhao1,2 Yiming Ju1,2 Chengwei Wu1 Tengfei Pan
@@
{ldu, hyzhao, ymju, cwwu, tfpan}@baai.ac.cn
@@ 2 Introduction
Instruction tuning serves as a cornerstone for unleashing the vast knowledge and reasoning capabilities of large pretrained models (Longpre et al., 2023; Ouyang et al., 2022). Consequently, the construction of high-quality instruction datasets has become essential for improving model performance and generalization ability. Several efforts have constructed instruction datasets through manual annotation or automatic synthesis, with the total size of these datasets having reached tens of millions (Ahmad et al., 2025; Nayak et al., 2024). However, current large models still struggle with complex instruction-following tasks and show limitation in “rare” tasks with low frequency (Qin et al., 2025).
A main reason would be that existing works overlooked the distribution of existing data, and thus failed to efficiently enhance the “coverage and “depth of synthesized instructions, and thus limit the model performance in solving rare tasks and difficult tasks (Wang et al., 2024; Liu et al., 2024; Long et al., 2024). The “coverage of an instruction dataset refers to the coverage of task types and knowledge categories it covers, while the ‘depth reflects the complexity of instructions within individual tasks. Both two dimensions are essential to enlarge the capability boundaries of large models (Shengyu et al., 2023). Expanding the coverage improves the model’s ability to generalize across domains, while increasing depth helps enhance complex reasoning ability and compositional generalizability (Sinha et al., 2024; Zhong et al., 2025). However, a key challenge remains: how to design effective, controllable and interpretable instruction generation strategies that continuously expand the instruction space along both dimensions, guiding models to acquire higher-level capabilities (Zhang et al., 2023; Qian et al., 2022; Qin et al., 2022). Currently, there is a lack of unified methodology and empirical research in this area.
@@
Based on this framework, we construct the Infinity Instruct Subject (InfInstruct-Sub) dataset, which contains $\sim$1.5 million high-quality instructions. The code and dataset are available at https://github.com/BAAI-DIPL/InfinityInstruct-Sub and https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen. Evaluations on multiple benchmarks demonstrate its effectiveness in improving instruction-following ability. Across foundation models, our tuned models outperform their official instruction-tuned counterparts. Data analysis further reveals that, compared with comparable synthetic datasets such as Magpie (Wei et al., 2023) and UltraChat (Ding et al., 2023), our framework could obtain an instruction set with higher coverage and complexity. Moreover, interestingly, we observe a scaling law in the co-occurrence structure of tags of instructions: the co-occurrence frequency of a tag with others follows a negative logarithmic relationship with its own frequency. This indicates a scale-free connection structure between instructions, similar to the topology of the Internet. Such a pattern may offer new insights into the scaling laws of model performance and suggest ways to better understand the internal knowledge structure of datasets, potentially improving training efficiency and effectiveness. This further highlights the importance of understanding the content structure of instruction data, and underscores the value of the high-quality tagging system proposed in this work.
@@ 3 Construction of Infinity Instruct Subject
The construction process of Infinity Instruct Subject (InfInstruct-Sub) is illustrated in Figure 1. The process starts by collecting high-quality seed instructions. We systematically gathered all general-domain instruction datasets both manually created and automatically generated using GPT-4(Achiam et al., 2023) or ChatGPT available up to March 2024, resulting in a data pool of approximately 7 million samples. The included datasets are listed in Table 1 of the Appendix. Next, we developed an automatic tagging system based on large language models (LLMs) to analyze the distribution of the data pool. From the perspectives of coverage and depth, we selected a set of high-information seed instructions. Then, we applied an evolutionary algorithm to generate over one million new instruction samples by evolving towards greater complexity and difficulty. Building on both the seed and synthesized data, we constructed a model deficiency diagnosis system. This system identifies gaps in model capabilities and guides the targeted synthesis of new data to efficiently address those weaknesses. Finally, we implemented a strict semantic similarity–based data leakage prevention framework to detect and mitigate potential risks of data leakage throughout the construction process.
@@ 3.1 Hierarchical Multilingual Tagging System for Characterizing Instruction Content Distribution
Since LLMs may describe the same knowledge or skill using different expressions (e.g., “math calculation” vs. “mathematical calculation”), we normalize tags based on semantic similarity. Specifically, we obtain embeddings for all tags using BGE (Chen et al., 2024), and identify semantically similar tags whose cosine similarity exceeds an empirical threshold of $\lambda=0.91$. Within each group of similar tags, we retain the one with the highest frequency as the representative. To further refine the tag set, we apply DBSCAN (Ester et al., 1996) clustering with parameters $eps=0.47$ and min_samples=2min\_samples=2 to merge closely related tags. These thresholds are chosen based on empirical observations. After normalization and clustering, tags with frequency less than 100 are considered noisy and removed, following prior work (Lu et al., ). As a result, a total of 21,378 fine-grained tags are retained.
@@
Given the set of fine-grained tags, we select the top 1,000 most frequent ones, then use GPT-4 to automatically summarize them into a set of domain-level categories, so that we employ each induced category as a domain-level tag. To establish a mapping between fine-grained tags and domain-level tags, we use the prompt shown in Figure 4 along with the categorization criteria illustrated in Figure 1 of the Appendix. Based on this setup, we employ Qwen2.5-72B-Instruct (Yang et al., 2024) to generate the mappings between fine-grained and domain tags.
@@ 4. Undertrained Instructions:
In particular, we use Llama-2-7B-base (Touvron et al., 2023) to compute the loss values for instructions. To estimate the reduction ratio in token-level average cross-entropy loss, we fine-tune Llama-2-7B-base on a balanced instruction subset with 1,000 instructions per domain category. Table 1 presents the sources of the seed data and the number of instructions retained after filtering. By applying the above criteria sequentially, we select approximately 1.2 million informative instructions from the overall data pool.
@@ 3.3 Evolutionary Algorithm for In-Distribution Generalization
1. Metadata-Guided Evolution: In this step, we adopt the widely-adopted Evol-Instruct (Xu et al., 2023) method. By randomly choose one of four dimensions—diversity, i.e., more reasoning steps, concretizing, deepening, and more diversified, to guide each instruction’s evolution using State-of-the-Art advanced large model as synthesizer.
@@
3. Multi-Turn Dialogue Generation: Following (Ding et al., 2023), each valid instruction is used to generate 1–4 rounds of dialogue, with the AI assistant simulating different roles.
@@ 3.5 Data Leakage Prevention
Overlap between training and evaluation data can cause models to achieve inflated scores on familiar samples, leading to overestimated performance. We observe that many public instruction datasets include questions from common benchmarks like MT-Bench (Zheng et al., 2023b) and AlpacaEval (Li et al., 2023), which risks misleading evaluation. To address this issue, we use the BGE model to compute semantic similarity between evaluation queries and instructions in the training data. This process involves multi-dimensional comparisons, including semantic understanding and structural features, to ensure accurate and comprehensive matching. Based on the similarity scores, we filter out overlapping entries, effectively preventing data leakage and preserving the reliability of model evaluation.
@@ 4.1 Experimental Setup
We finetune open-source pretrained models Qwen2-7B-base and LLaMA3-8B-base on the InfInstruct-Sub dataset, and compare their performance against their respective official instruction-tuned and alignment-tuned versions. Evaluations are conducted on widely adopted LLM-judging-based benchmarks AlpacaEval 2.0 (Li et al., 2023) and Arena-Hard-V0.1 (Li et al., 2024). The reference model for AlpacaEval 2.0 is GPT-4-1106; on Arena-Hard-V0.1 is GPT-4-0314.
@@ 4.2 Experimental Results
+Self-Instruct (Wang et al., 2023) 100k 7.21 4.0
+ShareGPT (Zheng et al., 2023b) 112k 9.73 6.5
+Evol Instruct (Xu et al., 2023) 143k 8.52 5.1
+OpenHermes 1 (Teknium, 2023a) 243K 9.94 4.4
+Tulu V2 Mix (Ivison et al., 2023) 326K 9.91 5.4
+WildChat (Zhao et al., ) 652k 14.62 8.7
+OpenHermes 2.5 (Teknium, 2023b) 1M 12.89 8.2
+GenQA (University of Maryland, 2024) 6.47M 9.05 3.0
+UltraChat (Ding et al., 2023) 208k 8.29 3.6
+Magipie (Xu et al., 2024) 300k 22.66 14.9
@@ 5.1 Dataset Distribution
We further compare the semantic coverage of InfInstruct-Sub with similar instruction datasets, including Alpaca(Taori et al., 2023), llm-sys(Zheng et al., 2023a), and Magpie (Xu et al., 2024). We sample 20,000 instructions from each dataset and apply the embedding and projection method same to abovementioned. As shown in Figure 12, InfInstruct-Sub exhibits a broader semantic spread, demonstrating that our strategy of selecting more low-frequency instructions as seeds effectively increases coverage in the synthesized data.
@@
To evaluate the difficulty of synthesized data, we follow the approach proposed in Magpie (Xu et al., 2024), using a large language model to assign difficulty scores to instruction samples. The difficulty is rated on a five-point scale: ’very easy’: 0, ’easy’: 1, ’medium’: 2, ’hard’: 3, ’very hard’: 4. This evaluation process is independent of the method we use for selecting high-difficulty seed instructions. Specifically, we employ Qwen-2.5-32B-Instruct as the scorer and randomly sample 1,500 instructions from each of the following datasets: Alpaca-GPT, LLM-Sys, Magpie, and InfInstruct-Sub.
@@ 5.3 Scaling Phenomenon in Label Connectivity Distribution
The presence of this topology highlights the interdependent structure of knowledge in instruction data, which should be considered within the data construction, instruction selection, and model training process. This also highlights the necessity of understanding the distribution and interrelation of instruction contents (Zhao et al., 2025), and the necessity of a high-quality tagging system, which is the foundation of understanding such knowledge structure.
Moreover, prior work (e.g., (Tan et al., 2024)) hypothesizes that the scaling law between model performance and dataset size is originated from the distribution of “latent skills” within the data, if such “latent skills” follow a power-law distribution. Our findings provide empirical support for this hypothesis, offering new insights into the origin of scaling laws between data distribution and model performance. We leave further exploration of this direction to future work, which further emphasizes the importance of the tagging system introduced in this paper.
@@ 6.1 Instruction Data Synthesis
Manually constructed datasets rely on experts to write instructions and responses, such as LIMA (Zhou et al., 2023) and Dolly (Ding et al., 2023). While these datasets are high in quality, they are expensive to scale due to the high cost of human annotation.
@@
Semi-automatic approaches, like Self-Instruct(Wang et al., 2023), Alpaca(Taori et al., 2023), and Evol-Instruct(Muennighoff et al., 2023), expand a small set of human-labeled data using prompt engineering. While they improve scalability, the reliance on handcrafted prompts limits diversity and complexity. Fully automatic methods (e.g., WebInstruct(Yue et al., 2024), backtranslation) extract instruction-like data from web documents with minimal human input, but often lack precise control over coverage and difficulty. Recent studies explore improvements from three angles:
@@
  • Seed Selection and High-Information Filtering: The choice of seed data largely determines the downstream diversity and quality. Traditional methods often use a few high-frequency, simple instructions, leading to limited coverage(Wang et al., 2023). Recent work addresses this by leveraging hierarchical tag systems and multi-dimensional metrics to select high-information seeds—i.e., rare, diverse, and complex instructions that better represent undercovered regions of the task space(Tu et al., 2024; He et al., 2025; Xie et al., 2023). This ensures that synthesized data can expand both the coverage (domain/task diversity) and depth (instruction complexity) of the dataset.
@@
  • Evolution-Based Instruction Generation: To address the need for more challenging instructions, evolution-based methods iteratively expand seed data using strategies like genetic algorithms (Xu et al., 2023) or model feedback loops(Gu et al., 2025). These approaches increase instruction complexity and reasoning depth, while dynamically adjusting generation based on model performance. They also help avoid redundancy by diversifying synthesis paths.
@@
  • Instruction Synthesis Strategies: Most existing synthesis methods rely on prompting large language models with handcrafted or auto-generated templates(Wang et al., 2023; Xu et al., 2023). While effective, this can lead to templated and repetitive outputs. Magpie(Xu et al., 2024) proposes prompt-free instruction synthesis via autoregressive alignment, which improves fluency and semantic richness. Similarly, Web Reconstruction (Jiang et al., 2025) leverages real-world text to generate instruction-style data with higher realism and contextual diversity.
@@
Despite these advances, many datasets still underrepresent complex, multi-step, or domain-specific instructions. Few existing works explicitly address the coverage and depth of instruction distribution in a unified manner. This limits model generalization in scenarios requiring deep reasoning, long-range context, or adaptation to unfamiliar domains (Yu et al., 2024; Zhang et al., 2025). Moreover, while some studies explore feedback-driven data augmentation or generative instruction synthesis (Le et al., 2022), most lack a closed-loop framework that jointly optimizes coverage, difficulty, and model adaptation. The interplay between seed selection, model feedback, and evolution remains underexplored.